if the area if a triangular sail is 10 square feet and the base is 4 feet long, find the sail height
Answer by dolly(163)

Let the height be h ft
Base = 4 ft
So area = 1/2 * b * h
= 1/2 * 4 * h
= 2h
But given area = 10
==> 2h = 10
==> h = 10/2
==> h = 5
Thus the ht = 5 ft
Good luck!!!